OpenStudy (anonymous):

how high up on the building will a 15ft. ladder reach if the foot of the ladder is placed set from the building

OpenStudy (lukebluefive):

How far away from the building is the ladder's base?

OpenStudy (anonymous):

how high up on the building will a 15ft. ladder reach if the foot of the ladder is placed 5ft from the building

OpenStudy (anonymous):

sorry set is 5ft

OpenStudy (lukebluefive):

Using the Pythagorean Theorem, given two sides of a right triangle, you can find the length of the other side: \[side_1^2 + side_2^2 = hypotenuse^2\] In this case, we have: \[5^2 + side^2 = 15^2\] Rewritten: \[side^2 = 15^2 - 5^2\] Subtracting: \[side^2 = 200\] Taking the square root of both sides: \[side = \sqrt{200}\]